On March 31st, 2017 Alexander Corporation sold equipment for cash of $25,000. The equipment had been purchased on June 1st, 2015 for $42,000 and at that time they had estimated a $15,000 residual value at the end of the equipment's 4-year useful life. Alexander Corporation has a December 31st year end, adjusts its records annually, and uses the double declining/diminishing balance method of depreciation for this equipment.

Prepare the appropriate journal entrie(s) to record the disposal of the equipment on
March 31st, 2017. Please show all calculations to receive full marks. Round all amounts to the nearest dollar.
Answer #1

Double declining depreciation rate = 2 x 1/useful life

= 2 x 1/4

= 50%

depreciation expense for year 2015 = 42,000 x 50% x 7/12

= $12,250

Book value of equipment at January 1, 2016 = Cost price - Depreciation expense for 2015

= 42,000-12,250

= $29,750

Depreciation expense for year 2016 = 29,750 x 50%

= $14,875

Book value of equipment at January 1, 2017 = Book value of equipment at January 1, 2016 - Depreciation expense for year 2016

= 29,750-14,875

= $14,875

Depreciation expense for year 2017 = 14,875 x 50% x 3/12

= $1,859

Accumulated depreciation till March 31, 2017 = 12,250+14,875+1,859

= $28,984

Book value of equipment on March 31, 2017 = Cost price - Accumulated depreciation

= 42,000-28,984

= $13,016

Sale price of equipment = $25,000

Gain on sale of equipment = Sale price of equipment- Book value of equipment on March 31, 2017

= 25,000-13,016

= $11,984

Date General Journal Debit Credit
March 31,2017 Cash 25,000
Accumulated depreciation- equipment 28,984
Equipment 42,000
Gain on disposal of equipment 11,984

Kindly comment if you need further assistance. Thanks‼!